# TEARS: Textual Representations for Scrutable Recommendations

### v1

@@ 1. Introduction
Figure 3. Controllability experiment visualization: large-scope changes (top/grey frame), fine-grained edits (middle/orange frame), and guided recommendations (bottom/ purple frame). Red indicates edited summaries, green are base summaries and blue are models. Summaries and examples are paraphrased. App. P includes more summaries, with examples in App. G for large-scope and App. H for fine-grained.
@@ 2. Related Work
Scrutable recommender systems, in contrast, have been limited in flexibility, largely relying on keyword or tag-based methods (Gao et al., 2023; Gedikli et al., 2014; Green et al., 2009; Vig et al., 2009; Balog et al., 2019; Mysore et al., 2023), where users can toggle tags on or off. However, this approach can hinder controllability, as users are burdened with parsing through numerous keywords or tags, leading to significant cognitive load. Few works have explored scrutable systems through natural language user summaries. Sanner et al. (2023) conducted a user study that highlighted the advantages of using user summaries to generate zero or few-shot recommendations with a frozen LLM. The study found that under specific conditions, LLMs leveraging user summaries can compete with black-box models in cold-start settings. Closest to TEARS, Ramos et al. (2024) developed a scrutable model using natural language profiles for next-item prediction. However, the approach relies on user reviews, which are not always available, and only approaches the performance of black-box models. In contrast, our work does not require user reviews and employs a hybrid methodology to demonstrate that TEARS can surpass black-box model performance and maintain scrutability.
@@ 3.2. Summary Design
While LLMs have shown impressive capabilities in text summarization (El-Kassas et al., 2021), we find that free-style prompting without adherence to a specific structure can make summaries generic and vary in quality. On the other hand, LLMs have excelled in instruction-based tasks (Ouyang et al., 2022). With this in mind, we design a prompt asking for summaries to include the desirable characteristics mentioned to enforce consistency and quality. Consistent summaries will also likely help train a decoder and obtain high-quality recommendations. Our resulting prompting strategy is in Figure 8. We explicitly direct the model to avoid stating titles and rating information to prevent over-reliance on such details and encourage summaries to be more expressive. By design, LLM’s responses are non-deterministic. In early experiments, we find that summary generation can vary, with two output summaries being significantly different for the same user. We observe higher variability in users with longer histories, leading us to limit the number of items used for each summary (to a maximum of 50 items in our studies). We further explore this effect in App. M. Note that the number of items used for each summary, $m_{u}$, is user-dependent since some users have shorter histories than this max. threshold. Finally, our prompt also contains the expected length of the summaries, which we set to 200 words, as we find it short enough not to incur heavy cognitive loads but can be detailed enough. We leave further exploration on the impact of summary length and if it should vary by user for future work.
@@ 3.3. TEARS
where $\lambda_{1}$ and $\lambda_{2}$ are weighing parameters for their respective losses. In practice, we initialize $D$ with the base model’s decoder, update its weights while training, and freeze $Q_{r}$’s weights (see App. L.2).
@@ 5.1. User Summary Properties
We analyze the average summary length to inspect if summaries can be appropriately scrutinized. For GPT, summaries average $168.6{\scriptsize\pm 3.6}$ words across all datasets, compared to $179.6{\scriptsize\pm 2.2}$ for LlaMA. These lengths suggest that the summaries are concise enough to be easily editable while still comprehensive enough to convey detailed user information, as reflected in their positive impact on recommendation performance (see §5.2). To assess uniqueness, we use pairwise edit distance and BLEU scores (Papineni et al., 2001), the latter measuring n-gram overlap between two texts (we use 4-grams). The average edit distance for GPT summaries is $160.1{\scriptsize\pm 0.1}$ words, while LlaMA summaries average $157.0{\scriptsize\pm 1.2}$. These scores are high when compared to the average summary length, indicating distinct summaries across users. Similarly, BLEU scores are low, with GPT summaries averaging $0.08{\scriptsize\pm 0.02}$ and LlaMA summaries $0.19{\scriptsize\pm 0.01}$, which suggest minimal n-gram overlap, so diverse phrasing. Further details and statistics are in App. A and examples in App. P.
@@ 5.2. Recommendation Quality
Overall, TEARS VAE models show that it is possible to significantly enhance scrutability without sacrificing—and even improving—performance compared to both AE and genre-based models. TEARS provides a plug-in to standard AEs. We provide additional analysis over components of TEARS, such as ablation studies in App. L, analysis on user activity levels in App. N, and evaluate GPT-based models using LLaMA summaries and vice versa in App. O.
Figure 4. Tradeoff between recommendation performance (y-axis) and large scope controllability (x-axis) for ML-1M (left) and Goodbooks (middle) using GPT-generated summaries (see App. F for LLaMA results). Netflix data is shown in Figure 11. The x-axis represents $|\Delta_{\text{up/down}}|$ as $\alpha$ increases, reflecting its impact on NDCG@20. Notably, most TEARS VAE models outperform TEARS BASE in recommendation performance at $\alpha=1$, with some also achieving superior controllability. The bar plots (right) illustrate guided recommendation outcomes, where all models successfully guide black-box embeddings in the intended direction. Results are averaged across five seeds.All details are in App. K.
@@ 7. Ablation Studies
We perform ablation studies to compare different pooling methodologies for TEARS and different values of the hyperparameter responsible for the OT procedure $\lambda_{1}$ (see Equation 6). We perform these on the ML-1M dataset to assess the efficacy of the proposed method. All methods are trained using the same random seed and the hyperparameters from the best-performing TEARS MacridVAE model, using GPT-4-turbo generated summaries. We provide further ablations in App L, where we analyze the various components of the loss function (App. L.1), what weights should be trained and which should be frozen (App. L.2) and pre-initializing the text-encoder for TEARS VAEs using a trained TEARS Base model (App. L.3).
@@ Appendix I Genre Based Controllability
Appendix L Ablations
@@
L.1. Loss Function Configurations
@@
L.2. What Weights to Train
@@
L.3. Using TEARS Base to Initialize the Text-Encoder
@@
Appendix M Stochasticity in GPT Generated Summaries
@@
Appendix N Cold Start Experiment
@@
Appendix O Generalization to Different LLMs
@@
Appendix P GPT-4-preview Example Summaries
P.1. ML-1M
@@
Appendix Q LLaMA 3.1 Example Summaries
Q.1. ML-1M
@@
Q.4. Summary Generation Prompts

### v2

@@ 1. Introduction
Figure 3. Controllability experiment visualization: large-scope changes (top/grey frame), fine-grained edits (middle/orange frame), and guided recommendations (bottom/ purple frame). Red indicates edited summaries, green are base summaries and blue are models. Summaries and examples are paraphrased. App. Q includes more summaries, with examples in App. G for large-scope and App. H for fine-grained.
@@ 2. Related Work
Scrutable recommender systems, in contrast, have been limited in flexibility, largely relying on keyword or tag-based methods (Gao et al., 2023; Gedikli et al., 2014; Green et al., 2009; Vig et al., 2009; Balog et al., 2019; Mysore et al., 2023), where users can toggle tags on or off. However, this approach can hinder controllability, as users are burdened with parsing through numerous keywords or tags, leading to significant cognitive load. Few works have explored scrutable systems through natural language user summaries. Sanner et al. (2023) conducted a user study that highlighted the advantages of using user summaries to generate zero or few-shot recommendations with a frozen LLM. The study found that under specific conditions, LLMs leveraging user summaries can compete with black-box models in cold-start settings. Moreover, Ramos et al. (2024) developed a scrutable model using natural language profiles for next-item prediction. However, the approach relies on user reviews, which are not always available, and only approaches the performance of black-box models. In contrast, our work does not require user reviews and employs a hybrid methodology to demonstrate that TEARS can surpass black-box model performance and maintain scrutability. Additionally, there has been interest in studying the role of user profiles for scrutable recommendations by some concurrent work, specifically analyzing qualitative properties such as length or readability (Zhou et al., 2024) as well as how to train a language model to enhance the quality of user summaries using a RS as a reward function (Gao et al., 2025).
@@ 3.2. Summary Design
While LLMs have shown impressive capabilities in text summarization (El-Kassas et al., 2021), we find that free-style prompting without adherence to a specific structure can make summaries generic and vary in quality. On the other hand, LLMs have excelled in instruction-based tasks (Ouyang et al., 2022). With this in mind, we design a prompt asking for summaries to include the desirable characteristics mentioned to enforce consistency and quality. Consistent summaries will also likely help train a decoder and obtain high-quality recommendations. Our resulting prompting strategy is in Figure 8. We explicitly direct the model to avoid stating titles and rating information to prevent over-reliance on such details and encourage summaries to be more expressive. By design, LLM’s responses are non-deterministic. In early experiments, we find that summary generation can vary, with two output summaries being significantly different for the same user. We observe higher variability in users with longer histories, leading us to limit the number of items used for each summary (to a maximum of 50 items in our studies). We further explore this effect in App. N. Note that the number of items used for each summary, $m_{u}$, is user-dependent since some users have shorter histories than this max. threshold. Finally, our prompt also contains the expected length of the summaries, which we set to 200 words, as we find it short enough not to incur heavy cognitive loads but can be detailed enough. We leave further exploration on the impact of summary length and if it should vary by user for future work.
@@ 3.3. TEARS
where $\lambda_{1}$ and $\lambda_{2}$ are weighing parameters for their respective losses. In practice, we initialize $D$ with the base model’s decoder, update its weights while training, and freeze $Q_{r}$’s weights (see App. M.2). Additionally, in App. L we compare OT to other common alignment methodologies such as JS-divergence (symmetric KL) and contrastive learning finding that OT is superior to both techniques in our setting.
@@ 5.1. User Summary Properties
We analyze the average summary length to inspect if summaries can be appropriately scrutinized. For GPT, summaries average $168.6{\scriptsize\pm 3.6}$ words across all datasets, compared to $179.6{\scriptsize\pm 2.2}$ for LlaMA. These lengths suggest that the summaries are concise enough to be easily editable while still comprehensive enough to convey detailed user information, as reflected in their positive impact on recommendation performance (see §5.2). To assess uniqueness, we use pairwise edit distance and BLEU scores (Papineni et al., 2001), the latter measuring n-gram overlap between two texts (we use 4-grams). The average edit distance for GPT summaries is $160.1{\scriptsize\pm 0.1}$ words, while LlaMA summaries average $157.0{\scriptsize\pm 1.2}$. These scores are high when compared to the average summary length, indicating distinct summaries across users. Similarly, BLEU scores are low, with GPT summaries averaging $0.08{\scriptsize\pm 0.02}$ and LlaMA summaries $0.19{\scriptsize\pm 0.01}$, which suggest minimal n-gram overlap, so diverse phrasing. Further details and statistics are in App. A and examples in App. Q.
@@ 5.2. Recommendation Quality
Overall, TEARS VAE models show that it is possible to significantly enhance scrutability without sacrificing—and even improving—performance compared to both AE and genre-based models. TEARS provides a plug-in to standard AEs. We provide additional analysis over components of TEARS, such as ablation studies in App. M, analysis on user activity levels in App. O, and evaluate GPT-based models using LLaMA summaries and vice versa in App. P.
[…]
@@ 7. Ablation Studies
We perform ablation studies to compare different pooling methodologies for TEARS and different values of the hyperparameter responsible for the OT procedure $\lambda_{1}$ (see Equation 6). We perform these on the ML-1M dataset to assess the efficacy of the proposed method. All methods are trained using the same random seed and the hyperparameters from the best-performing TEARS MacridVAE model, using GPT-4-turbo generated summaries. We provide further ablations in App M, where we analyze the various components of the loss function (App. M.1), what weights should be trained and which should be frozen (App. M.2) and pre-initializing the text-encoder for TEARS VAEs using a trained TEARS Base model (App. M.3).
@@ Appendix I Genre Based Controllability
Appendix L Comparison With Other Alignment Methods
In this work we primarily explored using an OT alignment due the imposed normal prior on the VAE models we are aling to. In this Appendix, we focus on comparing against some other alignment strategies and providing some theoretical insights as to why OT works the best.
Table 10. Scores for varying alignment strategies. We overall find that OT works best for our setting yielding the best recommendation performance and controllability.
NDCG@50c NDCG@50s $|\Delta_{\text{down}}@20|$ $|\Delta_{\text{up}}@20|$
No alignment ($\lambda_{1}=0$) 0.291 0.251 0.261 0.410
OT 0.296 0.273 0.470 0.740
JSD 0.291 0.257 0.271 0.249
Contrastive 0.293 0.256 0.378 0.582
To further demonstrate the importance of the OT objective, we compare OT against two popular alignment strategies: Jensen-Shannon Divergence (JSD) and contrastive alignment. For this comparison, we tune $\lambda_{1}$ for each method and report the run achieving the highest average NDCG@50 over $\alpha=\{0,0.5,1\}$, which is equivalent to the OT setting.
For JSD, we utilize its closed-form solution when evaluated using normal priors:
$D_{JS}(P\|Q)$ $=\frac{1}{2}D_{KL}(P\|M)+\frac{1}{2}D_{KL}(Q\|M),$
where $M=\frac{1}{2}(P+Q)$.
$D_{KL}(P\|Q)$ $=\int p(x)\log\frac{p(x)}{q(x)}dx,$
$D_{KL}(P\|Q)$ $=\log\frac{\sigma_{Q}}{\sigma_{P}}+\frac{\sigma_{P}^{2}+(\mu_{P}-\mu_{Q})^{2}}{2\sigma_{Q}^{2}}-\frac{1}{2}.$
For contrastive learning, we use the contrastive loss presented by Ren et al. (2024):
(15) $\mathcal{L}_{CL}=-\mathbb{E}\log\left(\frac{f(s_{i},e_{i})}{\sum_{j}f(s_{j},e_{i})}\right),$
where $s_{i}=\text{concat}[\mu_{s,i},\sigma_{s,i}]$ and $e_{i}=\text{concat}[\mu_{r,i},\sigma_{r,i}]$, with $\sigma_{s/r}$ representing the diagonal elements of the covariance matrix. The function $f(s_{i},e_{i})$ is defined as:
$f(s_{i},e_{i})=\exp\left(\frac{s_{i}e_{i}^{T}}{\|s_{i}\|_{2}\|e_{i}\|_{2}}\right).$
Table 10 presents the scores for all alignment methods. The results indicate that OT consistently outperforms the alternatives in both controllability and recommendation performance. Interestingly, we observe that using JSD for alignment offers minimal improvement compared to using no alignment at all. We hypothesize that this is because JSD serves as a suboptimal optimization target relative to OT, as noted in Arjovsky et al. (2017). When the compared distributions exhibit little to no overlap, JSD becomes highly sensitive to small changes, complicating the optimization process. Additionally, OT demonstrates superior performance over contrastive learning. We attribute this to limitations in contrastive learning’s ability to fully align the Gaussian parameters between the two distributions. Specifically, in our setup, $f(s_{i},e_{i})$ is the only term in $\mathcal{L}_{CL}$ that aligns the Gaussian parameters for user $i$. The optimum of $\log f(s_{i},e_{i})=\frac{s_{i}e_{i}^{T}}{\|s_{i}\|_{2}\|e_{i}\|_{2}}$ is invariant to the norms of $s_{i}$ and $e_{i}$, meaning contrastive learning aligns only the directions of the Gaussian parameters, not their magnitudes. In contrast, OT explicitly incorporates both direction and magnitude into its optimization target (Equation 1), which likely explains its superior performance. We note that contrastive learning may be advantageous in scenarios where the model’s latents do not follow parameterized distributions, offering greater flexibility in such cases. However, exploring these settings is beyond the scope of this work and is left for future research.
Appendix M Ablations
@@
M.1. Loss Function Configurations
@@
M.2. What Weights to Train
@@
M.3. Using TEARS Base to Initialize the Text-Encoder
@@
Appendix N Stochasticity in GPT Generated Summaries
@@
Appendix O Cold Start Experiment
@@
O.1. Additional Decoder LM
In this work, we explore using a base-T5 model. To validate that our using a base T5 model is not a strict requirement, we perform an ablation by substituting it with a base BERT model. Table 15 displays the scores for BERT when compared against the T5 models. While we find T5 outperforms BERT, we note that it still outperforms the standalone RecVAE. Additionally, we find that similar to T5, OT helps improve the controllability and recommendation performance of the model.
NDCG@50c NDCG@50s $|\Delta_{\text{down}}@20|$ $|\Delta_{\text{up}}@20|$
T5 w OT 0.296 0.273 0.470 0.740
BERT w OT 0.282 0.240 0.267 0.398
BERT w.o OT 0.278 0.231 0.209 0.2598
MacridVAE 0.276 N/A N/A N/A
Appendix P Generalization to Different LLMs
@@
Appendix Q GPT-4-preview Example Summaries
Q.1. ML-1M
@@
Appendix R LLaMA 3.1 Example Summaries
R.1. ML-1M
@@
R.4. Summary Generation Prompts